# Value-oriented forecast reconciliation for renewables in electricity markets

### v1

@@ 3.2 Value-oriented Forecast Reconciliation
Our reconciliation model is inspired by Pennings and Van Dalen, (2017). Specifically, the reconciliation model $f_{\texttt{RE}}$ is implemented using an aggregation matrix $\mathbf{S}$ combined with a learnable combination function $g$. Additionally, we incorporate contextual features, such as market data, represented as $\mathbf{z}_{t}\in\mathbb{R}^{d}$, throughout the reconciliation process. As a result, both the base forecasts $\hat{\mathbf{y}}_{t+k|t}$ and the contextual features $\mathbf{z}_{t}$ serve as inputs to the model $f_{\texttt{RE}}$. The reconciliation procedure is characterized by
@@
where $\mathbf{h}_{t+k}$ denotes the latent variable and $\bm{\theta}$ symbolizes the parameters within the combination function $g$. In accordance with (12), $\mathbf{h}_{t+k}$ can also be interpreted as the bottom-level forecasts. More specifically, the learning-based reconciliation process uses $g$ to generate reconciled bottom-level forecasts, which are subsequently aggregated to obtain forecasts at the aggregate level.
@@
For example, Ben Taieb and Koo, (2019) use the widely adopted mean squared error as the loss function, aiming to minimize the total root mean square error. In contrast, here each WPP $i$ has the cost function $c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$ and seeks to minimize its own cost. Therefore, for a group of WPPs, while each aims to minimize its own costs, they must collectively decide on the parameters $\bm{\theta}$ within the reconciliation model. Under the classical utilitarian principle, the parameter estimation procedure involves minimizing the sum of costs for all WPPs, expressed as:
$\min_{\bm{\theta}}\sum_{i=1}^{m}\mathbb{E}\left[c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})\right].$
The parameters obtained from this estimation are referred to as the utilitarian estimate. However, the sum of costs is neutral with respect to potential inequalities in the distribution of costs among the involved agents. In contrast, incorporating fairness considerations into the parameter estimation process offers a viable alternative. To achieve this, we use the cooperative bargaining framework to distribute costs fairly among the agents involved.
@@ 3.2.1 Nash Bargaining based Reconciliation
$l_{i}(\bm{\theta})=\left\{\begin{aligned} \mathbb{E}\left[c_{t+k,i}^{(\texttt{AG})}(f_{\texttt{RE}}(\hat{\mathbf{y}}_{t+k|t},\mathbf{z}_{t};\bm{\theta}),\mathbf{y}_{t+k})\right],&\ \texttt{if}\ \bm{\theta}\in\mathcal{A},\\ \mathbb{E}\left[c_{t+k,i}(\hat{y}_{t+k,i},y_{t+k,i})\right],&\ \texttt{if}\ \bm{\theta}=\mathbf{d}.\end{aligned}\right.$ (20)
This suggests that when involved agents reach a consensus, they will reconcile their forecasts and trade collectively in the markets, thereby sharing the associated costs. If they fail to reach a consensus, they will trade independently. The disagreement point $\mathbf{d}$ can be assigned an unattainable value, such as an all-zero vector, since the reconciliation of forecasts is no longer required. Specifically, we assume that within the agreement set, there are points that present Pareto improvements [^1] Consider a set of functions $l_{1},l_{2},\cdots,l_{m}$ to be minimized using parameters $\bm{\theta}$. $\bm{\theta}$ is a Pareto improvement to $\bm{\theta}^{\prime}\in\mathcal{A}$, if $l_{i}(\bm{\theta})\leq l_{i}(\bm{\theta}^{\prime}),\forall i=1,2\cdots,m$. over the disagreement point. If no such points exist, individual rationality will be violated, thereby rendering the design of a reconciliation function meaningless. This assumption is intuitively plausible, as Theorem 2.1 demonstrates how aggregate trading can reduce overall balancing costs. We will show that it holds within the framework of our allocation rule under a few mild conditions.
Particularly, the Nash bargaining framework proposes four axioms, namely Pareto optimality, symmetry, independence of irrelevant alternatives, and invariance to affine transformations (Nash,, 1953), which are described as follows:
@@
    Pareto optimality: the solution $\bm{\theta}^{\texttt{NB}}$ is Pareto optimal in $\mathcal{A}$, that is, there does not exist $\bm{\theta}\in\mathcal{A}$ such that $l_{i}(\bm{\theta})<l_{i}(\bm{\theta}^{\texttt{NB}}),\ \forall i=1,2,\cdots,m$, and $\bm{\theta}\neq\bm{\theta}^{\texttt{NB}}$.
@@
    Symmetry: the solution $\bm{\theta}^{\texttt{NB}}$ is invariant in the order of agents.
@@
    Independence of irrelevant alternatives: if we expand the feasible set $\mathcal{A}$ and the solution stays in $\mathcal{A}$, then the solution $\bm{\theta}^{\texttt{NB}}$ stays the same.
@@
According to Nash, (1953), a unique solution exists for the bargaining game. In the context of the cost sharing game defined in this study, the Nash bargaining solution can be derived as follows:
@@
$\bm{\theta}^{\texttt{NB}}=\arg\underset{\bm{\theta}\in\mathcal{A}}{\min}$ $\sum_{i}^{m}-\log\left[l_{i}(\mathbf{d})-l_{i}(\bm{\theta})\right]$ (21a)
s.t. $l_{i}(\mathbf{d})\geq l_{i}(\bm{\theta}),\ i=1,2,\cdots,m.$ (21b)
3.2.2 The Derivation of Cost Allocation
@@
We design the cost allocation rule $c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$ for each WPP within the framework of value-oriented forecast reconciliation. Specifically, we employ the weighted proportional allocation rule introduced by Anupindi et al., (2001). This rule consists of two components: one associated with the bottom-level reconciled forecast $\tilde{y}_{t+k,i|t}$ and the other with the aggregate-level reconciled forecast $\tilde{y}_{t+k,\texttt{sum}|t}$. The cost $c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$ is mathematically expressed as
$c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})\,=\,(1-w)\,c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i})\,+\,w\,\gamma_{i}\,c_{t+k}(\tilde{y}_{t+k,\texttt{sum}|t},y_{t+k,\texttt{sum}}).$ (22)
@@
where $0\leq w\leq 1$, $\gamma_{i}\geq 0$, and $\sum_{i}^{m}\gamma_{i}=1$. The term $c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i})$ represents the cost associated with the WPP $i$ if it independently submits the pseudo-offer $\tilde{y}_{t+k,i|t}$ in the markets. In contrast, $c_{t+k}(\tilde{y}_{t+k,\texttt{sum}|t},y_{t+k,\texttt{sum}})$ represents the total cost for the aggregation when the PM submits the offer $\tilde{y}_{t+k,\texttt{sum}|t}$ in the markets.
@@
This allocation rule implies that each WPP reduces a fraction of its own cost associated with the pseudo-offer while simultaneously bearing an equivalent fraction of its share of the overall system cost, $\gamma_{i}c_{t+k}(\tilde{y}_{t+k,\texttt{sum}|t},y_{t+k,\texttt{sum}})$. When $w=0$, the PM retains the entire profit from the aggregate trading and distributes none to the WPPs. Consequently, each WPP acts independently using the reconciled forecast $\tilde{y}_{t+k,i|t}$, as if aggregation does not exist. In contrast, when $w=1$, the cost incurred by each WPP is represented by $\gamma_{i}c_{t+k}(\tilde{y}_{t+k,\texttt{sum}|t},y_{t+k,\texttt{sum}})$, which denotes a portion of the system-wide cost. For $w\in(0,1)$, the weighted proportional allocation rule reflects allocation schemes that fall between these extreme cases.
@@
Specifically, $\gamma_{i}$ can be defined according to various criteria, as outlined by Hartman and Dror, (1996). For example, one definition uses the proportion of actual power generation, given by:
$\gamma_{i}^{\texttt{ge}}=\frac{y_{t+k,i}}{y_{t+k,\texttt{sum}}}.$
@@
$\gamma_{i}^{\texttt{pc}}=\frac{c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i})}{\sum_{i=1}^{m}c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i})}.$
@@
Designate $\gamma_{i}$ in the cost allocation formula (22) as $\gamma_{i}^{\texttt{pc}}$. Thus, we obtain the following inequality:
$c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})\leq c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i}),$
which ensures that the cost allocation for WPP $i$ is almost surely less than the corresponding cost associated with its pseudo offers.
@@
Define $\gamma_{i}$ in the cost allocation (22) as $\gamma_{i}^{\texttt{ge}}$. If the following condition holds:
$\frac{c_{t+k}(\tilde{y}_{t+k,\texttt{sum}|t},y_{t+k,\texttt{sum}})}{y_{t+k,\texttt{sum}}}\leq\frac{c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i})}{y_{t+k,i}},\ \forall i=1,2,\ldots,m,$ (23)
@@
$c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})\leq c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i}).$
The proof is straightforward and thus omitted here. In practice, the condition (23) is typically satisfied, as demonstrated through simulations and case studies. If the cost $c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i})$ for the pseudo offer $\tilde{y}_{t+k,i|t}$ is lower than that of the independent offer $c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})$, then the aforementioned propositions suggest that the cost experienced by WPP $i$ under the proposed method, represented by $c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$, is less than $c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})$. Nevertheless, it implies a reduction in the cost of independent offers post-reconciliation, which is a strong requirement.
Remark 3.1.
For the BU reconciliation approach, as the bottom-level forecasts remain unchanged post-reconciliation, i.e., $\tilde{y}_{t+k,i|t}=\hat{y}_{t+k,i|t}$, it is evident that $c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$ is less than $c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})$.
Remark 3.2.
For the proposed value-oriented forecast reconciliation approach, if we can find parameters in the agreement set, then $\mathbb{E}\left[c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})\right]$ is less than $\mathbb{E}\left[c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})\right]$, as given by the Nash bargaining framework (21).
It is noted that, using $\gamma_{i}^{\texttt{pc}}$ disrupts the convexity of $c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$, thereby complicating the loss function $l_{i}(\bm{\theta})$. Consequently, we mainly utilize $\gamma_{i}^{\texttt{ge}}$ in this work, as it preserves the convexity of $c_{t+k,i}^{(\texttt{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$. To further analyze the cost allocations, we formally define the extra profit from the forecast reconciliation as follows.

### v2

@@ 3.2 Value-oriented Forecast Reconciliation
Our reconciliation model is inspired by pennings2017integrated. Specifically, the reconciliation model $f_{\mathrm{RE}}$ is implemented using a structural matrix $\mathbf{S}$ combined with a learnable combination function $g$. Additionally, we incorporate contextual features, such as market data, represented as $\mathbf{z}_{t}\in\mathbb{R}^{d}$, throughout the reconciliation process. As a result, both the base forecasts $\hat{\mathbf{y}}_{t+k|t}$ and the contextual features $\mathbf{z}_{t}$ serve as inputs to the model $f_{\mathrm{RE}}$. The reconciliation procedure is characterized by
@@
where $\mathbf{h}_{t+k}$ denotes the latent variable and $\bm{\theta}$ symbolizes the parameters within the combination function $g$. In accordance with (12), $\mathbf{h}_{t+k}$ can also be interpreted as the reconciled bottom-level forecasts. This approach generalizes the widely used reconciliation technique based on linear combinations of base forecasts (hollyman2021understanding), by allowing the input vector $\mathbf{z}_{t}$ to incorporate information beyond the original time series. Specifically, the learning-based reconciliation process leverages both the time series and external information to produce reconciled forecasts at the bottom level, which are then aggregated to generate forecasts at higher levels. This strategy aligns with the framework proposed by bertani2024joint.
@@
For example, ben2019regularized use the widely adopted mean squared error as the loss function, aiming to minimize the total root mean square error. In contrast, here each WPP $i$ has the cost function $c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$ and seeks to minimize its own cost. Therefore, for a group of WPPs, they must collectively decide on the parameters $\bm{\theta}$ within the reconciliation model. In such a multi-agent system, the parameter estimation procedure, following the classical utilitarian principle, entails minimizing the total cost incurred by all WPPs, formally expressed as:
$\min_{\bm{\theta}}\sum_{i=1}^{m}\mathbb{E}\left[c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})\right].$
The parameters obtained from this estimation are referred to as the utilitarian estimate. However, while each WPP aims to minimize its own costs, the sum of costs is neutral with respect to potential inequalities in the distribution of costs among the involved agents. Although weighting factors can be incorporated into the reconciliation process to account for forecasts at each hierarchical level (van2015game), there is a lack of well-defined methodologies for determining these weights in the context of decision-making. In contrast, incorporating fairness considerations into the parameter estimation process offers a viable and rigorous alternative (bertsimas2011price). To achieve this, we use the cooperative bargaining framework to distribute costs fairly among the agents involved.
@@ 3.2.1 Nash Bargaining based Reconciliation
$l_{i}(\bm{\theta})=\left\{\begin{aligned} \mathbb{E}\left[c_{t+k,i}^{(\mathrm{AG})}(f_{\mathrm{RE}}(\hat{\mathbf{y}}_{t+k|t},\mathbf{z}_{t};\bm{\theta}),\mathbf{y}_{t+k})\right],&\ \mathrm{if}\ \bm{\theta}\in\mathcal{A},\\ \mathbb{E}\left[c_{t+k,i}(\hat{y}_{t+k,i},y_{t+k,i})\right],&\ \mathrm{if}\ \bm{\theta}=\mathbf{d}.\end{aligned}\right$. (22)
This implies that when the participating agents reach a consensus, they reconcile their forecasts and engage in collective market trading, thereby sharing the associated costs. Under Assumption 3.1, sub-coalitions are not permitted; thus, in the absence of consensus, each agent trades independently. The disagreement point $\mathbf{d}$ can be assigned an unattainable value, such as an all-zero vector, since the reconciliation of forecasts is no longer required. Specifically, we assume that within the agreement set, there are points that present Pareto improvements [^1] Consider a set of objective functions $l_{1},l_{2},\cdots,l_{m}$ to be minimized with respect to a parameter vector $\bm{\theta}$. We say that $\bm{\theta}$ is a Pareto improvement over $\bm{\theta}^{\prime}\in\mathcal{A}$, if $l_{i}(\bm{\theta})\leq l_{i}(\bm{\theta}^{\prime}),\forall i=1,2\cdots,m$, and the inequality is strict for at least one index $i$. over the disagreement point. If no such points exist, individual rationality will be violated, thereby rendering the design of a reconciliation function meaningless. This assumption is intuitively plausible, as Theorem 2.1 demonstrates how aggregate trading can reduce overall balancing costs. We will show that it holds within the framework of our allocation rule under a few mild conditions.
Particularly, the Nash bargaining framework proposes four axioms, namely Pareto optimality, symmetry, independence of irrelevant alternatives, and invariance to affine transformations (nash1953two), which are described as follows:
@@
    Pareto optimality: the solution $\bm{\theta}^{\mathrm{NB}}$ is Pareto optimal in $\mathcal{A}$, that is, there does not exist $\bm{\theta}\in\mathcal{A}$ such that $l_{i}(\bm{\theta})<l_{i}(\bm{\theta}^{\mathrm{NB}}),\ \forall i=1,2,\cdots,m$, and $\bm{\theta}\neq\bm{\theta}^{\mathrm{NB}}$.
@@
    Symmetry: the solution $\bm{\theta}^{\mathrm{NB}}$ is invariant in the order of agents.
@@
    Independence of irrelevant alternatives: if we expand the feasible set $\mathcal{A}$ and the solution stays in $\mathcal{A}$, then the solution $\bm{\theta}^{\mathrm{NB}}$ stays the same.
@@
According to nash1953two, a unique solution exists for the bargaining game. In the context of the cost sharing game defined in this study, the Nash bargaining solution can be derived as follows:
$\bm{\theta}^{\mathrm{NB}}=\arg\underset{\bm{\theta}\in\mathcal{A}}{\min}\$$\sum_{i}^{m}-\log\left[l_{i}(\mathbf{d})-l_{i}(\bm{\theta})\right]$ (23a)
$\mathrm{s.t.}\$$l_{i}(\mathbf{d})\geq l_{i}(\bm{\theta}),\ i=1,2,\cdots,m.$ (23b)
The constraint in (23) ensures that the allocated cost for each agent does not exceed the cost incurred under independent offering, as stipulated by Assumption 3.2.
In fact, neither $\mathbb{E}\left[c_{t+k,i}^{(\mathrm{AG})}(f_{\mathrm{RE}}(\hat{\mathbf{y}}_{t+k|t},\mathbf{z}_{t};\bm{\theta}),\mathbf{y}_{t+k})\right]$ nor $\mathbb{E}\left[c_{t+k}(\hat{y}_{t+k,i},y_{t+k,i})\right]$ is available. Following ban2019big, we employ the sample average approximation method to approximate the expectation components in (23) and apply the empirical risk minimization algorithm for parameter estimation. As a result, we convert problem (23) into a constrained learning problem. With a set of historical data $\{\hat{\mathbf{y}}_{t+k|t},\mathbf{z}_{t},\mathbf{y}_{t+k}\}_{t=1,2,\cdots,T}$, we can rewrite (23) as
@@
$\underset{\bm{\theta}}{\min}$ $\,\quad\sum_{i}^{m}-\log\left[\frac{1}{T}\sum_{t=1}^{T}\left[c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})-c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})\right]\right],$ (24a)
$\quad\mathrm{s.t.}$ $\quad\,\frac{1}{T}\sum_{t=1}^{T}\left[c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})-c_{t+k}(\hat{y}_{t+k,i},y_{t+k,i})\right]\leq 0,\ \forall i=1,\cdots,m,$ (24b)
Unlike ban2019big, which employs linear models for data-driven decisions, we leverage neural networks to model the combination function $g$, as described in (19). In the following subsections, we describe the allocation rule used and defer the details of model training to Section 3.3.
3.2.2 Weighted Proportional Allocation Mechanism
@@
We design the cost allocation rule $c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$ for each WPP within the framework of value-oriented forecast reconciliation. Specifically, we employ the weighted proportional allocation rule introduced by anupindi2001general. This rule consists of two components: one associated with the bottom-level reconciled forecast $\tilde{y}_{t+k,i|t}$ and the other with the aggregate-level reconciled forecast $\tilde{y}_{t+k,\mathrm{sum}|t}$. The cost $c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$ is mathematically expressed as
$c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})\,=\,(1-w)\,c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i})\,+\,w\,\gamma_{i}\,c_{t+k}(\tilde{y}_{t+k,\mathrm{sum}|t},y_{t+k,\mathrm{sum}}).$ (25)
@@
where $0\leq w\leq 1$, $\gamma_{i}\geq 0$, and $\sum_{i}^{m}\gamma_{i}=1$. The term $c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i})$ represents the cost associated with the WPP $i$ if it independently submits the pseudo-offer $\tilde{y}_{t+k,i|t}$ in the markets. In contrast, $c_{t+k}(\tilde{y}_{t+k,\mathrm{sum}|t},y_{t+k,\mathrm{sum}})$ represents the total cost for the aggregation when the PM submits the offer $\tilde{y}_{t+k,\mathrm{sum}|t}$ in the markets.
@@
This allocation rule implies that each WPP reduces a fraction of its own cost associated with the pseudo-offer while simultaneously bearing an equivalent fraction of its share of the overall system cost, $\gamma_{i}c_{t+k}(\tilde{y}_{t+k,\mathrm{sum}|t},y_{t+k,\mathrm{sum}})$. When $w=0$, the PM retains the entire profit from the aggregate trading and distributes none to the WPPs. Consequently, each WPP acts independently using the reconciled forecast $\tilde{y}_{t+k,i|t}$, as if aggregation does not exist. In contrast, when $w=1$, the cost incurred by each WPP is represented by $\gamma_{i}c_{t+k}(\tilde{y}_{t+k,\mathrm{sum}|t},y_{t+k,\mathrm{sum}})$, which denotes a portion of the system-wide cost. For $w\in(0,1)$, the weighted proportional allocation rule reflects allocation schemes that fall between these extreme cases.
@@
Specifically, $\gamma_{i}$ can be defined according to various criteria, as outlined by hartman1996cost. For example, one definition uses the proportion of actual power generation [^2] When $y_{t+k,\mathrm{sum}}$ is equal to 0, we set $\gamma_{i}^{\mathrm{ge}}$ as $1/m$., given by:
$\gamma_{i}^{\mathrm{ge}}=\frac{y_{t+k,i}}{y_{t+k,\mathrm{sum}}}.$
@@
$\gamma_{i}^{\mathrm{pc}}=\frac{c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i})}{\sum_{i=1}^{m}c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i})}.$
@@
Designate $\gamma_{i}$ in the cost allocation formula (25) as $\gamma_{i}^{\mathrm{pc}}$. Thus, we obtain the following inequality:
$c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})\leq c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i}),$
which ensures that the cost allocation for WPP $i$ is less than or equal to the corresponding cost associated with its pseudo offers.
@@
Define $\gamma_{i}$ in the cost allocation (25) as $\gamma_{i}^{\mathrm{ge}}$. If the following condition holds:
$\frac{c_{t+k}(\tilde{y}_{t+k,\mathrm{sum}|t},y_{t+k,\mathrm{sum}})}{y_{t+k,\mathrm{sum}}}\leq\frac{c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i})}{y_{t+k,i}},\ \forall i=1,2,\ldots,m,$ (26)
@@
$c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})\leq c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i}).$
The proof is straightforward and thus omitted here. In practice, the condition (26) is typically satisfied, as demonstrated through simulations and case studies. If the cost $c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i})$ for the pseudo offer $\tilde{y}_{t+k,i|t}$ is lower than that of the independent offer $c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})$, then the aforementioned propositions suggest that the cost experienced by WPP $i$ under the proposed method, represented by $c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$, is less than $c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})$. Nevertheless, it implies a reduction in the cost of independent offers post-reconciliation, which is a strong requirement.
We note that Propositions 3.2 and 3.3 hold for any parameters $\bm{\theta}$. Particularly, in the context of the BU reconciliation approach, the bottom-level forecasts remain unchanged after reconciliation, i.e., $\tilde{y}_{t+k,i|t}=\hat{y}_{t+k,i|t}$. Consequently, it follows that the shared allocation cost $c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$ is less than or equal to the cost of independent offering $c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})$. It is evident that,
$\mathbb{E}\left[c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})\right]\leq\mathbb{E}\left[c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})\right].$
Based on this result, we derive the following corollary.
Corollary 3.2.
By employing the weighted proportional allocation rule, the agreement set of parameters $\mathbf{\theta}$ in problem (23) is guaranteed to be non-empty.
Given that both $c_{t+k}(\tilde{y}_{t+k,i|t},y_{t+k,i})$ and $c_{t+k}(\tilde{y}_{t+k,\mathrm{sum}|t},y_{t+k,\mathrm{sum}})$ are convex functions, introducing $\gamma_{i}^{\mathrm{pc}}$ disrupts the convexity of the cost function $c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$, as the ratio of convex functions is generally not convex. In contrast, $\gamma_{i}^{\mathrm{ge}}$ is a constant and thus preserves the convexity of $c_{t+k,i}^{(\mathrm{AG})}(\tilde{\mathbf{y}}_{t+k|t},\mathbf{y}_{t+k})$. For this reason, we primarily adopt $\gamma_{i}^{\mathrm{ge}}$ throughout this work.
3.2.3 Profit Allocation within the Aggregation
To further analyze the profit allocations, we formally define the extra profit from the forecast reconciliation as follows.